# PI3DETR: Parametric Instance Detection of 3D Point Cloud Edges with a Geometry-Aware 3DETR

### v1

@@ Abstract
We present PI3DETR, an end-to-end framework that directly predicts 3D parametric curve instances from raw point clouds, avoiding the intermediate representations and multi-stage processing common in prior work. Extending 3DETR, our model introduces a geometry-aware matching strategy and specialized loss functions that enable unified detection of differently parameterized curve types, including cubic Bézier curves, line segments, circles, and arcs, in a single forward pass. Optional post-processing steps further refine predictions without adding complexity. This streamlined design improves robustness to noise and varying sampling densities, addressing critical challenges in real world LiDAR and 3D sensing scenarios. PI3DETR sets a new state-of-the-art on the ABC dataset and generalizes effectively to real sensor data, offering a simple yet powerful solution for 3D edge and curve estimation.
@@ 2 Related work
As noted by Zhu et al. [25], a common strategy for parametric curve reconstruction is to first detect geometric primitives such as corners and edges and then then fit curves in a separate post processing stage [23, 6, 13, 11]. PIE-NET [23] follows this paradigm using a PointNet++ [16] backbone to classify points into corner, edge, or background classes, after which a secondary network predicts curve proposals between detected corner pairs and estimates their parameters. The method supports lines, circles, and B-splines, and is trained on the ABC dataset [9].
@@ 3 Method
3.1 Dataset and ground truth.
@@ 3.2 PI3DETR Overview
For Bézier curves, $\boldsymbol{\theta}^{\mathbf{B}}$ maps $o^{\text{f}}_{j}$ to four control points $\hat{\mathbf{B}}_{j}=[\hat{\mathbf{b}}^{\text{s}}_{j},\hat{\mathbf{b}}^{\text{c1}}_{j},\hat{\mathbf{b}}^{\text{c2}}_{j},\hat{\mathbf{b}}^{\text{e}}_{j}]\in\mathbb{R}^{4\times 3}$. For line segments, $\boldsymbol{\theta}^{\mathbf{L}}$ outputs the midpoint $\hat{\mathbf{l}}^{\text{m}}_{j}\in\mathbb{R}^{3}$ and unit direction $\hat{\mathbf{l}}^{\text{v}}_{j}\in\mathbb{R}^{3}$, while $\boldsymbol{\theta}^{\ell}$ predicts the length $\hat{l}^{\text{r}}_{j}\in\mathbb{R}$. For circles, $\boldsymbol{\theta}^{\mathbf{C}}$ returns the center $\hat{\mathbf{c}}^{\text{m}}_{j}\in\mathbb{R}^{3}$ and unit normal $\hat{\mathbf{c}}^{\text{v}}_{j}\in\mathbb{R}^{3}$, and $\boldsymbol{\theta}^{r}$ estimates the radius $\hat{c}^{\text{r}}_{j}\in\mathbb{R}$. We summarize the line segment and circle predictions as $\hat{\mathbf{L}}_{j}=[\hat{\mathbf{l}}^{\text{m}}_{j},\hat{\mathbf{l}}^{\text{v}}_{j},\hat{l}^{\text{r}}_{j}]$ and $\hat{\mathbf{C}}_{j}=[\hat{\mathbf{c}}^{\text{m}}_{j},\hat{\mathbf{c}}^{\text{v}}_{j},\hat{c}^{\text{r}}_{j}]$, respectively. For arcs, $\boldsymbol{\theta}^{\mathbf{A}}$ predicts the start point, mid-point, and end point $\hat{\mathbf{A}}_{j}=[\hat{\mathbf{a}}^{\text{s}}_{j},\hat{\mathbf{a}}^{\text{m}}_{j},\hat{\mathbf{a}}^{\text{e}}_{j}]\in\mathbb{R}^{3\times 3}$.
A classification head $\boldsymbol{\theta}^{\text{cls}}$ maps $o^{\text{f}}_{j}$ to $\hat{p}_{j}\in[0,1]^{5}$ over the set {no-object, cubic Bézier, line segment, circle, arc}, selecting the curve type at inference. PI3DETR computes parameters for all curve types from every $o^{\text{f}}_{j}$, while a geometry-aware matcher assigns gradients only to the heads corresponding to the ground-truth type during training.
@@
where $\mathcal{L}_{\text{CE}}(\hat{p}_{j},c_{i})$ is the cross-entropy loss between the predicted class probabilities $\hat{p}_{j}$ and ground truth class $c_{i}$, $\mathcal{L}_{\text{param}}$ is the geometry-aware parametric curve loss described earlier, and $\mathcal{L}_{\text{CD}}$ is the Chamfer distance (CD) loss. Here, $S^{c_{i}}_{j}$ and $S^{c_{i}}_{i}$ denote the $64$ uniformly sampled 3D points from the curves of class $c_{i}$ using the respective parameters of $j$ and $i$.
Predictions matched to a placeholder ground truth instance with $c_{i}=0$ contribute only to the class prediction branch through the cross entropy loss. We apply auxiliary losses after each decoder layer as in [3, 13] and since our dataset exhibits a class imbalance among curve types, we use class weights in the cross entropy loss. The computation of these weights and details of the CD loss are provided in Sec. B.2 and for completeness, CD is defined in Appendix C.
@@ 4.2 Stress testing
To assess robustness, we evaluate different point cloud densities $N$ by randomly sub-sampling the test set (Tab. 2). PI3DETR achieves the lowest CD and HD across all settings, with performance only degrading at N=4,096N=4{,}096. Our method outperforms all NerVE variants with N=8,192N=8{,}192 points for CD and N=4,096N=4{,}096 points for HD. In the sparsest case, NerVE CAD yields $347$ invalid outputs, whereas our method always produces valid predictions.
@@
To assess robustness against noise, we define three levels $\eta\in{\tfrac{s}{1e^{3}},\tfrac{s}{5e^{2}},\tfrac{s}{2e^{2}}}$, where $\eta$ is the standard deviation of Gaussian noise added to the 3D coordinates and $s$ denotes the largest axis-aligned span of the point cloud. Results in Tab. 3 show that PI3DETR consistently outperforms NerVE CAD across all noise levels for both CD and HD, and also surpasses NerVE PWL on HD. For CD, NerVE PWL performs better at lower noise levels, but PI3DETR achieves the best score once noise increases to $\eta=\tfrac{s}{2e^{2}}$. Although PI3DETR degrades relative to the noise-free setting, it remains consistently more robust than NerVE as noise increases. The stronger CD of NerVE PWL at low noise can be attributed to spurious PWLs detected in noisy point clouds, which lower CD but increase HD. This effect is visible in Fig. 7, which shows predictions of all methods for $\eta=\tfrac{s}{2e^{2}}$, and is consistent with the higher HD values reported in Tab. 3 or an invalid prediction percentage of 46%.
@@ 5 Conclusion
We presented PI3DETR, a streamlined end-to-end framework for detecting differently parameterized 3D curve instances in point clouds in a single forward pass. By incorporating a geometry-aware matching strategy, PI3DETR removes the need for intermediate representations or multi-stage processing and can optionally use simple post-processing steps. Our method achieves state-of-the-art performance, lowering CD on the ABC dataset by almost half compared to NerVE, and is particularly robust on sparse point clouds. PI3DETR’s limitations are the need for extensive training to achieve optimal results, the lack of a fallback for misclassified curves, and the challenge of directly regressing curves without explicit point associations. Nevertheless, its simplicity makes it a powerful foundation for future advances in 3D curve detection and geometric reasoning, paving the way toward more structured and intelligent 3D understanding.
@@ B.1 Geometry-aware matching
B.2 Loss
Class Weight Computation
@@ B.2 Loss
Chamfer Loss Computation

### v2

@@ Abstract
We present PI3DETR, an end-to-end framework that directly predicts 3D parametric curve instances from raw point clouds, avoiding the intermediate representations and multi-stage processing common in prior work. Extending 3DETR, our model introduces a geometry-aware matching strategy and specialized loss functions that enable unified detection of differently parameterized curve types, including cubic Bézier curves, line segments, circles, and arcs, in a single forward pass. Optional post-processing steps further refine predictions without adding complexity. This streamlined design improves robustness to noise and varying sampling densities, addressing critical challenges in real world LiDAR and 3D sensing scenarios. PI3DETR sets a new state-of-the-art on the ABC dataset and generalizes effectively to real sensor data, offering a simple yet powerful solution for 3D edge and curve estimation. Code: https://github.com/fafraob/pi3detr.
@@ 2 Related Work
As noted by Zhu et al. [25], a common strategy for parametric curve reconstruction is to first detect geometric primitives such as corners and edges and then then fit curves in a separate post-processing stage [23, 6, 13, 11]. PIE-NET [23] follows this paradigm using a PointNet++ [16] backbone to classify points into corner, edge, or background classes, after which a secondary network predicts curve proposals between detected corner pairs and estimates their parameters. The method supports lines, circles, and B-splines, and is trained on the ABC dataset [9].
@@ 3 Method
3.1 Dataset and Ground Truth
@@ 3.2 PI3DETR Overview
For Bézier curves, $\boldsymbol{\theta}^{\mathbf{B}}$ maps $o^{\text{f}}_{j}$ to four control points $\hat{\mathbf{B}}_{j}=[\hat{\mathbf{b}}^{\text{s}}_{j},\hat{\mathbf{b}}^{\text{c1}}_{j},\hat{\mathbf{b}}^{\text{c2}}_{j},\hat{\mathbf{b}}^{\text{e}}_{j}]\in\mathbb{R}^{4\times 3}$. For line segments, $\boldsymbol{\theta}^{\mathbf{L}}$ outputs the midpoint $\hat{\mathbf{l}}^{\text{m}}_{j}\in\mathbb{R}^{3}$ and unit direction $\hat{\mathbf{l}}^{\text{v}}_{j}\in\mathbb{R}^{3}$, while $\boldsymbol{\theta}^{\ell}$ predicts the length $\hat{l}^{\text{r}}_{j}\in\mathbb{R}$. For circles, $\boldsymbol{\theta}^{\mathbf{C}}$ returns the center $\hat{\mathbf{c}}^{\text{m}}_{j}\in\mathbb{R}^{3}$ and unit normal $\hat{\mathbf{c}}^{\text{v}}_{j}\in\mathbb{R}^{3}$, and $\boldsymbol{\theta}^{r}$ estimates the radius $\hat{c}^{\text{r}}_{j}\in\mathbb{R}$. We summarize the line segment and circle predictions as $\hat{\mathbf{L}}_{j}=[\hat{\mathbf{l}}^{\text{m}}_{j},\hat{\mathbf{l}}^{\text{v}}_{j},\hat{l}^{\text{r}}_{j}]$ and $\hat{\mathbf{C}}_{j}=[\hat{\mathbf{c}}^{\text{m}}_{j},\hat{\mathbf{c}}^{\text{v}}_{j},\hat{c}^{\text{r}}_{j}]$, respectively. For arcs, $\boldsymbol{\theta}^{\mathbf{A}}$ predicts the start point, mid-point, and end point $\hat{\mathbf{A}}_{j}=[\hat{\mathbf{a}}^{\text{s}}_{j},\hat{\mathbf{a}}^{\text{m}}_{j},\hat{\mathbf{a}}^{\text{e}}_{j}]\in\mathbb{R}^{3\times 3}$. A classification head $\boldsymbol{\theta}^{\text{cls}}$ maps $o^{\text{f}}_{j}$ to $\hat{p}_{j}\in[0,1]^{5}$ over the set {no-object, cubic Bézier, line segment, circle, arc}, selecting the curve type at inference. PI3DETR computes parameters for all curve types from every $o^{\text{f}}_{j}$, while a geometry-aware matcher assigns gradients only to the heads corresponding to the ground-truth type during training.
For simplicity, we describe all positions as if they were predicted directly. In practice, each point is obtained as an offset from its associated query location $q_{j}$, which stabilizes learning and improves localization accuracy.
@@
where $\mathcal{L}_{\text{CE}}(\hat{p}_{j},c_{i})$ is the cross-entropy loss between the predicted class probabilities $\hat{p}_{j}$ and ground truth class $c_{i}$, $\mathcal{L}_{\text{param}}$ is the geometry-aware parametric curve loss, and $\mathcal{L}_{\text{CD}}$ is the Chamfer distance (CD) loss. Here, $S^{c_{i}}_{j}$ and $S^{c_{i}}_{i}$ denote the $64$ uniformly sampled 3D points from the curves of class $c_{i}$ using the respective parameters of $j$ and $i$.
Predictions matched to a placeholder ground truth instance with $c_{i}=0$ contribute only to the class prediction branch through the cross entropy loss. We apply auxiliary losses after each decoder layer as in [3, 14] and since our dataset exhibits a class imbalance among curve types, we use class weights in the cross entropy loss. The weight computation and details of the CD loss are provided in Sec. B.3 and for completeness, CD is defined in Appendix C.
@@
We optionally apply two post-processing steps to refine PI3DETR predictions and suppress duplicates, with their impact analyzed in the ablation study (Sec. 4.3). Snap & Fit (S&F) snaps predicted curves to the point cloud and refits them using its predicted class (Fig. 4(a)), while an IoU Filter removes overlapping curves of the same class based on sampled-point IoU and confidence (Fig. 4(b)). More details are provided in Sec. B.2.
@@ 4.2 Stress Testing
To evaluate robustness, we vary point cloud density $N$ by randomly sub-sampling the test set (Tab. 2). PI3DETR consistently achieves the lowest CD and HD across all settings, with performance only degrading at N=4,096N=4{,}096. It surpasses all NerVE variants at N=8,192N=8{,}192 for CD and N=4,096N=4{,}096 for HD. In the sparsest setting, NerVE CAD produces 347 invalid outputs, while our method remains fully valid. To further validate our end-to-end design, we report results for the same sparse case without the S&F post-processing, achieving a CD of $0.0049$ and an HD of $0.0841$, surpassing even the full pipeline.
@@
To assess robustness against noise, we define three levels $\eta\in{\tfrac{s}{1e^{3}},\tfrac{s}{5e^{2}},\tfrac{s}{2e^{2}}}$, where $\eta$ is the standard deviation of Gaussian noise added to the 3D coordinates and $s$ denotes the largest axis-aligned span of the point cloud. Results in Tab. 3 show that PI3DETR consistently outperforms NerVE CAD across all noise levels for both CD and HD, and also surpasses NerVE PWL on HD. For CD, NerVE PWL performs better at lower noise levels, but PI3DETR achieves the best score once noise increases to $\eta=\tfrac{s}{2e^{2}}$. Without the post-processing step, PI3DETR attains a CD of $0.0138$ and an HD of $0.1986$, slightly weaker than the full pipeline yet still surpassing the NerVE baseline, further demonstrating its end-to-end robustness.
Although PI3DETR degrades relative to the noise-free setting, it remains consistently more robust than NerVE as noise increases. The stronger CD of NerVE PWL at low noise can be attributed to spurious PWLs detected in noisy point clouds, which lower CD but increase HD. This effect is visible in Fig. 7, which shows predictions of all methods for $\eta=\tfrac{s}{2e^{2}}$, and is consistent with the higher HD values reported in Tab. 3 or an invalid prediction percentage of 46%.
@@ 5 Conclusion
We presented PI3DETR, an end-to-end framework for detecting differently parameterized 3D curve instances in point clouds in a single forward pass. By incorporating a geometry-aware matching strategy, PI3DETR removes the need for intermediate representations or multi-stage processing and can optionally use simple post-processing steps. We achieve state-of-the-art performance, lowering CD on the ABC dataset by almost half compared to NerVE, and is particularly robust on sparse point clouds. PI3DETR’s limitations are the need for extensive training to achieve optimal results, the lack of a fallback for misclassified curves, and the challenge of directly regressing curves without explicit point associations. Nevertheless, its simplicity makes it a powerful foundation for future advances in 3D curve detection and geometric reasoning, paving the way toward more structured and intelligent 3D understanding.
@@ B.1 Geometry-Aware Matching
B.2 Optional Post-Processing
B.3 Loss
Class weight computation.
@@ B.3 Loss
Chamfer loss computation.